A man bought a goat for 60. He then sold it for 70. Later, he bought that goat back for 80, and then sold it again for 90. How much profit did he make?
1
Expert's answer
2020-09-09T11:07:18-0400

For the first part his profit was

Profit = selling price -buying price


"70-60= 10"

Profit was 10

The second part

Profit = selling price -buying price


"90-80=10"

The total profit therefore is


"10+10=20"